# Evan Munro    Xu Kuang    Stefan Wager

### v3

@@ B.2 Proof of Theorems 2, 3 and 4
$\mathbb{E}\left[\left(y_{n}\left(P_{n}(\bm{W})+U_{i})\right)-y_{n}\left(p^{*}_{\pi}\right)-\xi_{y}\xi_{z}^{-1}\frac{1}{n}\sum\limits_{i=1}^{n}Z_{i}(W_{i},p^{*}_{\pi}+U_{i})\right)^{2}\right]$
$\quad\quad\quad\quad=\mathbb{E}\left[\left(y\left(P_{n}(\bm{W})+U_{i})\right)-y\left(p^{*}_{\pi}\right)-\left(P_{n}(\bm{W})+U_{i})-p^{*}_{\pi}\right)y^{\prime}\left(p^{*}_{\pi}\right)\right)^{2}\right]+o(1/n)$
$\quad\quad\quad\quad=\mathcal{O}\left(\mathbb{E}\left[\left\lVert P_{n}(\bm{W})+U_{i})-p^{*}_{\pi}\right\rVert_{2}^{2}\right]\right)+o(1/n)=o(1/n),$
where on the last line we used continuous twice differentiability of $y(p)$ and the bound (43) to control the second moment of market price fluctuations. ∎
@@ B.3 Proof of Theorem 5
$\hat{\tau}_{\text{ADE}}-\tau^{*}_{\text{ADE}}=\frac{1}{n}\sum_{i=1}^{n}\frac{W_{i}\left(Y_{i}-y(1,\,p^{*}_{\pi})\right)}{\hat{\pi}}-\frac{(1-W_{i})\left(Y_{i}-y(1,\,p^{*}_{\pi})\right)}{1-\hat{\pi}}.$ (56)
@@
$\frac{1}{n}\sum_{i=1}^{n}\frac{W_{i}\left(Y_{i}-y(1,p^{*}_{\pi})\right)}{\pi}$ $=\frac{1}{n}\sum_{i=1}^{n}\frac{W_{i}\,\varepsilon_{i}(1)}{\pi}+y(1,\,P_{n})-y(1,\,p^{*}_{\pi})+o_{p}\left(1/\sqrt{n}\right)$
$=\frac{1}{n}\sum_{i=1}^{n}\frac{W_{i}\,\varepsilon_{i}(1)}{\pi}-\nabla_{p}^{\top}y(1,\,p^{*}_{\pi})\xi_{z}^{-1}Z_{i}(W_{i},\,p^{*}_{\pi})+o_{p}\left(1/\sqrt{n}\right),$
$\frac{1}{n}\sum_{i=1}^{n}\frac{(1-W_{i})\left(Y_{i}-y(0,p^{*}_{\pi})\right)}{1-\pi}$ $=\frac{1}{n}\sum_{i=1}^{n}\frac{(1-W_{i})\,\varepsilon_{i}(1)}{1-\pi}+y(0,\,P_{n})-y(0,\,p^{*}_{\pi})+o_{p}\left(1/\sqrt{n}\right)$
$=\frac{1}{n}\sum_{i=1}^{n}\frac{(1-W_{i})\,\varepsilon_{i}(1)}{1-\pi}-\nabla_{p}^{\top}y(0,\,p^{*}_{\pi})\xi_{z}^{-1}Z_{i}(W_{i},\,p^{*}_{\pi})+o_{p}\left(1/\sqrt{n}\right),$
and furthermore a central limit theorem follows by standard Lindeberg arguments,
$\frac{1}{\sqrt{n}}\begin{pmatrix}\sum_{i=1}^{n}\frac{W_{i}\left(Y_{i}-y(1,p^{*}_{\pi})\right)}{\pi}\\ \sum_{i=1}^{n}\frac{(1-W_{i})\left(Y_{i}-y(0,p^{*}_{\pi})\right)}{1-\pi}\end{pmatrix}\Rightarrow\left(\begin{pmatrix}0\\ 0\end{pmatrix},\ \operatorname{Cov}\left[\begin{pmatrix}\frac{W_{i}\,\varepsilon_{i}(1)}{\pi}-\nabla_{p}^{\top}y(1,\,p^{*}_{\pi})\xi_{z}^{-1}Z_{i}(W_{i},\,p^{*}_{\pi})\\ \frac{(1-W_{i})\,\varepsilon_{i}(0)}{1-\pi}-\nabla_{p}^{\top}y(0,\,p^{*}_{\pi})\xi_{z}^{-1}Z_{i}(W_{i},\,p^{*}_{\pi})\end{pmatrix}\right]\right).$
The result then follows from Slutsky’s lemma because $\hat{\pi}\rightarrow_{p}\pi$.
@@ B.4 Proof of Proposition 6
Throughout this proof, all quantities are computed in population equilibrium and so we use the short-hand $\Delta_{i}(w)=\Delta_{i}(w,\,p^{*}_{\pi})$. Expanding the square, we can verify that
$\sigma^{2}_{D}=\mathbb{E}\left[\frac{\varepsilon_{i}^{2}(1)}{\pi}+\frac{\varepsilon_{i}^{2}(0)}{1-\pi}\right]-2\mathbb{E}\left[\varepsilon_{i}(1)\Delta_{i}(1)-\varepsilon_{i}(0)\Delta_{i}(0)\right]+\mathbb{E}\left[\pi\Delta_{i}^{2}(1)+(1-\pi)\Delta_{i}^{2}(0)\right].$
Similarly, we have
$\bar{\sigma}^{2}_{D}$ $=\mathbb{E}\left[\left(\frac{1}{\pi}-1\right)\varepsilon_{i}^{2}(1)+\left(\frac{1}{1-\pi}-1\right)\varepsilon_{i}^{2}(0)+2\varepsilon_{i}(1)\varepsilon_{i}(0)\right]$
$\quad\quad-2\mathbb{E}\left[\left((1-\pi)\varepsilon_{i}(1)+\pi\varepsilon_{i}(0)\right)\left(\Delta_{i}(1)-\Delta_{i}(0)\right)\right]$
$\quad\quad+\mathbb{E}\left[\pi(1-\pi)\left(\Delta_{i}^{2}(1)-2\Delta_{i}(1)\Delta_{i}(0)+\Delta_{i}^{2}(0)\right)\right].$
Taking the difference and unexpanding squares where possible, we get
$\sigma^{2}_{D}-\bar{\sigma}^{2}_{D}$ $=\mathbb{E}\left[\left(\varepsilon_{i}(1)-\varepsilon_{i}(0)\right)^{2}\right]-2\mathbb{E}\left[\left(\varepsilon_{i}(1)-\varepsilon_{i}(0)\right)\left(\pi\Delta_{i}(1)+(1-\pi)\Delta_{i}(0)\right)\right]$
$\quad\quad\quad\quad\quad+\mathbb{E}\left[\left(\pi\Delta_{i}(1)+(1-\pi)\Delta_{i}(0)\right)^{2}\right]$
$=\mathbb{E}\left[\left(\varepsilon_{i}(1)-\varepsilon_{i}(0)-\left(\pi\Delta_{i}(1)+(1-\pi)\Delta_{i}(0)\right)\right)^{2}\right]$
$\geq 0,$
as claimed.
@@ B.5 Proof of Theorem 7
For simplicity, defined the unscaled perturbation $S_{i}=U_{i}/h_{n}$, and recall that $S_{i}\in\left\{\pm 1\right\}^{J}$ uniformly at random. Our estimator of the average indirect effect can equivalently be written as
$\hat{\tau}_{\text{AIE}}=-(\mathbf{Y}^{\top}\mathbf{S})\left(\mathbf{Z}^{\top}\mathbf{S}\right)^{-1}\hat{\tau}_{\text{ADE}}^{z}.$
One can replicate the proof of Theorem 5 to verify that $\hat{\tau}_{\text{ADE}}^{z}=z(1,\,p^{*}_{\pi})-z(0,\,p^{*}_{\pi})+\mathcal{O}_{P}(1/\sqrt{n})$, and thus makes an asymptotically negligible contribution to the error.
Next, by asymptotic equicontinuity as in Lemma 20,
$\frac{1}{n}\mathbf{Y}^{\top}\mathbf{S}$ $=\frac{1}{n}\sum_{i=1}^{n}Y_{i}(W_{i},\,p^{*}_{\pi}+h_{n}S_{i})S_{i}+\frac{h_{n}}{2^{J}}\sum_{v\in\left\{\pm 1\right\}^{J}}\left(y(P_{n}+h_{n}v)-y(p^{*}_{\pi}+h_{n}v)\right)v+o_{p}(1/\sqrt{n})$
$=\frac{1}{n}\sum_{i=1}^{n}Y_{i}(W_{i},\,p^{*}_{\pi}+h_{n}S_{i})S_{i}+\mathcal{O}(h_{n}^{2})+o_{p}(1/\sqrt{n})$
by twice continuous differentiability of $y$; a similar expansion holds for $\mathbf{Z}^{\top}\mathbf{S}$. Because $h_{n}=o(1/\sqrt[4]{n})$, we can then fold the $\mathcal{O}(h_{n}^{2})$ term into the $o_{p}(1/\sqrt{n})$ error. Again by twice differentiability of $y(\cdot)$, we can invoke Taylor expansion to verify that
$\text{vec}\left(Y_{i}(W_{i},\,p^{*}_{\pi}+h_{n}S_{i})\right)^{\top}\mathbf{S}\left(\mathbf{S}^{\top}\mathbf{S}\right)^{-1}$
@@
$\quad\quad\quad\quad\quad\quad+\text{vec}\left(Y_{i}(W_{i},\,p^{*}_{\pi}+h_{n}S_{i})-y(p^{*}_{\pi}+h_{n}S_{i})\right)^{\top}\mathbf{S}\left(\mathbf{S}^{\top}\mathbf{S}\right)^{-1}+o_{p}(1/\sqrt{n}),$
$\quad\quad=h_{n}\xi_{y}+\text{vec}\left(Y_{i}(W_{i},\,p^{*}_{\pi}+h_{n}S_{i})-y(p^{*}_{\pi}+h_{n}S_{i})+y(p^{*}_{\pi})\right)^{\top}\mathbf{S}\left(\mathbf{S}^{\top}\mathbf{S}\right)^{-1}+o_{p}(1/\sqrt{n}),$
and a similar decomposition holds for $\text{vec}\left(Z_{i}(W_{i},\,p^{*}_{\pi}+h_{n}S_{i})\right)^{\top}\mathbf{S}$. Furthermore
$\lim_{n\rightarrow\infty}\frac{1}{n}\sum_{i=1}^{n}\left(Y_{i}(W_{i},\,p^{*}_{\pi}+h_{n}S_{i})-y(p^{*}_{\pi}+h_{n}S_{i})-\left(Y_{i}(W_{i},\,p^{*}_{\pi})-y(p^{*}_{\pi})\right)\right)^{2}=0$
by weak continuity as in Assumptions 5 and 4, and these error terms are all mean-zero conditionally on $S_{i}$ and uncorrelated. Thus,
$\text{vec}\left(Y_{i}(W_{i},\,p^{*}_{\pi}+h_{n}S_{i})\right)^{\top}\mathbf{S}\left(\mathbf{S}^{\top}\mathbf{S}\right)^{-1}=h_{n}\xi^{\top}_{y}+\text{vec}\left(Y_{i}(W_{i},\,p^{*}_{\pi})\right)^{\top}\mathbf{S}\left(\mathbf{S}^{\top}\mathbf{S}\right)^{-1}+o_{p}(1/\sqrt{n});$
a similar expansion again holds for the $Z_{i}$ term.
@@
$(\mathbf{Y}^{\top}\mathbf{S})\left(\mathbf{Z}^{\top}\mathbf{S}\right)^{-1}$
$\quad=(\mathbf{Y}^{\top}\mathbf{S})\left(\mathbf{S}^{\top}\mathbf{S}\right)^{-1}\left(\mathbf{Z}^{\top}\mathbf{S}\left(\mathbf{S}^{\top}\mathbf{S}\right)^{-1}\right)^{-1}$
$\quad=\left(h_{n}\xi^{\top}_{y}+\mathbf{Y}^{*\top}\mathbf{S}\left(\mathbf{S}^{\top}\mathbf{S}\right)^{-1}\right)\left(h_{n}\xi_{z}+\mathbf{Z}^{*\top}\mathbf{S}\left(\mathbf{S}^{\top}\mathbf{S}\right)^{-1}\right)^{-1}+o_{p}\left(\frac{h_{n}}{\sqrt{n}}\right)$
$\quad=\xi^{\top}_{y}\xi_{z}^{-1}+h_{n}^{-1}\mathbf{Y}^{*\top}\mathbf{S}\left(\mathbf{S}^{\top}\mathbf{S}\right)^{-1}\xi_{z}^{-1}-h_{n}^{-1}\xi^{\top}_{y}\xi_{z}^{-1}\mathbf{Z}^{*\top}\mathbf{S}\left(\mathbf{S}^{\top}\mathbf{S}\right)^{-1}\xi_{z}^{-1}+o_{p}\left(\frac{h_{n}}{\sqrt{n}}\right),$
where the last step is by Taylor expansion. Finally, $n\left(\mathbf{S}^{\top}\mathbf{S}\right)^{-1}\xi_{z}^{-1}\rightarrow_{p}\xi_{z}^{-1}$, and so by Slutsky’s lemma
$\hat{\tau}_{\text{AIE}}-\tau^{*}_{\text{AIE}}=-\frac{1}{nh_{n}}\sum_{i=1}^{n}\left(Y_{i}(W_{i},\,p^{*}_{\pi})-\xi_{y}^{\top}\xi_{z}^{-1}Z_{i}(W_{i},\,p^{*}_{\pi})\right)S_{i}^{\top}\xi_{z}^{-1}\tau^{*}_{ADE,Z}+o_{p}\left(\frac{h_{n}}{\sqrt{n}}\right).$

### v4

@@ B.2 Proof of Theorems 2, 3 and 4
$\mathbb{E}\left[\left(\frac{1}{n}\sum\limits_{i=1}^{n}\left(y_{n}\left(P_{n}(\bm{W})\right)-y_{n}\left(p^{*}_{\pi}\right)\right)-\xi_{y}^{\top}\xi_{z}^{-1}\frac{1}{n}\sum\limits_{i=1}^{n}Z_{i}(W_{i},p^{*}_{\pi}+U_{i})\right)^{2}\right]$
$\quad\quad\quad\quad\leq\mathbb{E}\left[\left(\frac{1}{n}\sum\limits_{i=1}^{n}\left(y\left(P_{n}(\bm{W})\right)-y\left(p^{*}_{\pi}\right)-\xi_{y}^{\top}\left(P_{n}(\bm{W})-p^{*}_{\pi}\right)\right)\right)^{2}\right]+o(1/n)$
$\quad\quad\quad\quad=\mathcal{O}\left(\mathbb{E}\left[\left\lVert P_{n}(\bm{W})-p^{*}_{\pi}\right\rVert_{2}^{2}\right]\right)+o(1/n)$
$\quad\quad\quad\quad=o(1/n).$
On the second last line, we used the bound on the remainder term from Lemma 14. On the last line we used continuous twice differentiability of $y(p)$ and the bound (44) to control the second moment of market price fluctuations. ∎
@@ B.3 Proof of Theorem 5
$\hat{\tau}_{\text{ADE}}-\tau^{*}_{\text{ADE}}=\frac{1}{n}\sum_{i=1}^{n}\frac{W_{i}\left(Y_{i}-y_{n}(1,\,p^{*}_{\pi})\right)}{\hat{\pi}}-\frac{(1-W_{i})\left(Y_{i}-y_{n}(0,\,p^{*}_{\pi})\right)}{1-\hat{\pi}}.$ (60)
@@
$\frac{1}{n}\sum_{i=1}^{n}\frac{W_{i}\left(Y_{i}-y_{n}(1,p^{*}_{\pi})\right)}{\pi}$ $=\frac{1}{n}\sum_{i=1}^{n}\frac{W_{i}\,\varepsilon_{in}(1)}{\pi}+y_{n}(1,\,P_{n})-y_{n}(1,\,p^{*}_{\pi})+o_{p}\left(1/\sqrt{n}\right)$
$=\frac{1}{n}\sum_{i=1}^{n}\frac{W_{i}\,\varepsilon_{in}(1)}{\pi}-\nabla_{p}^{\top}y(1,\,p^{*}_{\pi})\xi_{z}^{-1}Z_{i}(W_{i},\,p^{*}_{\pi}+U_{i})+o_{p}\left(1/\sqrt{n}\right),$
$\frac{1}{n}\sum_{i=1}^{n}\frac{(1-W_{i})\left(Y_{i}-y_{n}(0,p^{*}_{\pi})\right)}{1-\pi}$ $=\frac{1}{n}\sum_{i=1}^{n}\frac{(1-W_{i})\,\varepsilon_{in}(0)}{1-\pi}+y_{n}(0,\,P_{n})-y_{n}(0,\,p^{*}_{\pi})+o_{p}\left(1/\sqrt{n}\right)$
$=\frac{1}{n}\sum_{i=1}^{n}\frac{(1-W_{i})\,\varepsilon_{in}(0)}{1-\pi}-\nabla_{p}^{\top}y(0,\,p^{*}_{\pi})\xi_{z}^{-1}Z_{i}(W_{i},\,p^{*}_{\pi}+U_{i})+o_{p}\left(1/\sqrt{n}\right),$
Then, taking the difference, we have
$\hat{\tau}_{\text{ADE}}-\tau^{*}_{\text{ADE}}=\frac{1}{n}\sum_{i=1}^{n}\frac{W_{i}\,\varepsilon_{in}(1)}{\pi}-\frac{(1-W_{i})\,\varepsilon_{in}(0)}{1-\pi}-\Delta_{in}(W_{i},p^{*}_{\pi})+o_{p}\left(1/\sqrt{n}\right),\$
where $\Delta_{in}(w,p)=\nabla_{p}^{\top}[y(1,p)-y(0,p)]\xi_{z}^{-1}Z_{i}(w,\,p+U_{i}).$ The central limit theorem follows by standard Lindeberg arguments and Slutsky’s theorem, since $\hat{\pi}\rightarrow_{p}\pi$. Note that the variance of the limiting distribution does not depend on $U_{i}$:
$\mathbb{E}\left[\left(\frac{W_{i}}{\pi}\varepsilon_{in}(1)+\frac{(1-W_{i})}{1-\pi}\varepsilon_{in}(0)-\Delta_{in}(W_{i},p^{*}_{\pi})\right)^{2}\right]=\mathbb{E}\left[\left(\frac{W_{i}}{\pi}\varepsilon_{i}(1)+\frac{(1-W_{i})}{1-\pi}\varepsilon_{i}(0)-\Delta_{i}(W_{i},p^{*}_{\pi})\right)^{2}\right]+o(1)$
by weak continuity in Assumption 6.
@@ B.4 Proof of Proposition 6
First, introduce the notation $a_{i}(1)=\epsilon_{i}(1)-\pi\Delta_{i}(1,p^{*}_{\pi})$, and $a_{i}(0)=\epsilon_{i}(0)+(1-\pi)\Delta_{i}(0,p^{*}_{\pi}).$ In an extension of the classical results in Neyman (1923), we can show that $\sigma^{2}_{D}$ is conservative for $\bar{\sigma}^{2}_{D}$.
$\begin{split}\sigma^{2}_{D}&=\mathbb{E}\left[\left(\frac{W_{i}\varepsilon_{i}(1)}{\pi}-\frac{(1-W_{i})\varepsilon_{i}(0)}{1-\pi}-\Delta_{i}(W_{i},p^{*}_{\pi})\right)^{2}\right],\\ &=\mathbb{E}\left[\left(\frac{W_{i}a_{i}(1)}{\pi}-\frac{(1-W_{i})a_{i}(0)}{1-\pi}\right)^{2}\right],\\ &=\mathbb{E}\left[\frac{a_{i}(1)^{2}}{\pi}\right]+\mathbb{E}\left[\frac{a_{i}(0)^{2}}{1-\pi}\right],\\ &=\mathbb{E}\left[\left(\frac{(1-\pi)a_{i}^{2}(1)}{\pi}+\frac{\pi a_{i}^{2}(0)}{1-\pi}+a^{2}_{i}(1)+a^{2}_{i}(0)\right)\right],\\ &\geq\mathbb{E}\left[\left(\frac{(1-\pi)a_{i}^{2}(1)}{\pi}+\frac{\pi a_{i}^{2}(0)}{1-\pi}+2\mathbb{E}[a_{i}(1)^{2}]^{1/2}\mathbb{E}[a_{i}(0)^{2}]^{1/2}\right)\right],\\ &\geq\mathbb{E}\left[\frac{(1-\pi)a_{i}^{2}(1)}{\pi}+\frac{\pi a_{i}^{2}(0)}{1-\pi}+2a_{i}(1)a_{i}(0)\right],\\ &=\pi\left(1-\pi\right)\mathbb{E}\left[\left(\frac{a_{i}(1)}{\pi}+\frac{a_{i}(0)}{1-\pi}\right)^{2}\right],\\ &=\pi\left(1-\pi\right)\mathbb{E}\left[\left(\frac{\varepsilon_{i}(1)}{\pi}+\frac{\varepsilon_{i}(0)}{1-\pi}-\left(\Delta_{i}(1,p^{*}_{\pi})-\Delta_{i}(0,p^{*}_{\pi})\right)\right)^{2}\right],\\ &=\bar{\sigma}_{D}^{2}\end{split}$
where the first inequality is from the AM-GM inequality and the second inequality is from Cauchy-Schwarz.
@@ B.5 Proof of Theorem 7
As usual we write unscaled perturbations as $S_{i}=U_{i}/h_{n}$, and recall that $S_{i}\in\left\{\pm 1\right\}^{J}$ uniformly at random. Our estimator of the average indirect effect can equivalently be written as
$\begin{split}&\hat{\tau}_{\text{AIE}}=-(\mathbf{Y}^{\top}\mathbf{S})\left(\mathbf{Z}^{\top}\mathbf{S}\right)^{-1}\hat{\tau}_{\text{ADE}}^{z}=-E_{Y}E_{Z}^{-1}\hat{\tau}_{\text{ADE}}^{z},\\ &E_{Y}=\mathbf{Y}^{\top}\mathbf{S}(\mathbf{S}^{\top}\mathbf{S})^{-1},\ \ \ \ E_{Z}=\mathbf{Z}^{\top}\mathbf{S}(\mathbf{S}^{\top}\mathbf{S})^{-1}.\end{split}$ (61)
One can replicate the proof of Theorem 5 to verify that $\hat{\tau}_{\text{ADE}}^{z}=z(1,\,p^{*}_{\pi})-z(0,\,p^{*}_{\pi})+\mathcal{O}_{P}(1/\sqrt{n})$. Meanwhile, as shown below, $E_{Y}$ and $E_{Z}$ will prove to have slower convergence rates. Thus, to capture the asymptotic behavior of $\hat{\tau}_{\text{AIE}}$ it will suffice to characterize $E_{Y}$ and $E_{Z}$; noise in $\hat{\tau}_{\text{ADE}}^{z}$ will make an asymptotically negligible contribution to the error.
We start by studying the term $E_{Y}$ below; the same argument can also be applied to $E_{Z}$. By asymptotic equicontinuity as in Lemma 22,
$\frac{1}{n}\mathbf{Y}^{\top}\mathbf{S}$ $=\frac{1}{n}\sum\limits_{i=1}^{n}Y_{i}(W_{i},p^{*}_{\pi}+h_{n}S_{i})S_{i}+(P_{n}(\bm{W})-p^{*}_{\pi})^{\top}\mathbb{E}[y^{\prime}(p^{*}_{\pi}+h_{n}S_{i})S_{i}]+o_{p}(1/\sqrt{n})$
$=\frac{1}{n}\sum\limits_{i=1}^{n}Y_{i}(W_{i},p^{*}_{\pi}+h_{n}S_{i})S_{i}+(P_{n}(\bm{W})-p^{*}_{\pi})^{\top}\mathbb{E}[S_{i}]y^{\prime}(p^{*}_{\pi})+O(h_{n}||P_{n}(\bm{W})-p^{*}_{\pi}||_{2})+o_{p}(1/\sqrt{n})$
$=\frac{1}{n}\sum_{i=1}^{n}Y_{i}(W_{i},\,p^{*}_{\pi}+h_{n}S_{i})S_{i}+o_{p}(1/\sqrt{n})$
The second step above is by a Taylor expansion of $y^{\prime}(p^{*}+h_{n}S_{i})$, using the assumption that the second derivatives of $y(p)$ are bounded and that $\mathbb{E}[S_{i}]=0$. Clearly $n^{-1}\mathbf{S}^{\top}\mathbf{S}\rightarrow_{p}I_{J\times J}$, and so the above implies that
$E_{Y}=\text{vec}\left(Y_{i}(W_{i},\,p^{*}_{\pi}+h_{n}S_{i})\right)^{\top}\mathbf{S}\left(\mathbf{S}^{\top}\mathbf{S}\right)^{-1}+o_{p}(1/\sqrt{n}).$
Furthermore, again by twice differentiability of $y(\cdot)$, we can invoke Taylor expansion to verify that
@@
$\quad+\text{vec}\left(Y_{i}(W_{i},\,p^{*}_{\pi}+h_{n}S_{i})-y(p^{*}_{\pi}+h_{n}S_{i})\right)^{\top}\mathbf{S}\left(\mathbf{S}^{\top}\mathbf{S}\right)^{-1}+o_{p}(1/\sqrt{n}),$
$=h_{n}\xi^{\top}_{y}+\text{vec}\left(Y_{i}(W_{i},\,p^{*}_{\pi})\right)^{\top}\mathbf{S}\left(\mathbf{S}^{\top}\mathbf{S}\right)^{-1}$
$\quad+\text{vec}\left(Y_{i}(W_{i},\,p^{*}_{\pi}+h_{n}S_{i})-y(p^{*}_{\pi}+h_{n}S_{i})-\left(Y_{i}(W_{i},\,p^{*}_{\pi})-y(p^{*}_{\pi})\right)\right)^{\top}\mathbf{S}\left(\mathbf{S}^{\top}\mathbf{S}\right)^{-1}+o_{p}(1/\sqrt{n}).$
Finally, by randomness of the $S_{i}$ and by weak continuity of net demand as given in Assumptions 6 (and as translated into outcomes by Assumption 5),
$\mathbb{E}\left[\left(Y_{i}(W_{i},\,p^{*}_{\pi}+h_{n}S_{i})-y(p^{*}_{\pi}+h_{n}S_{i})-\left(Y_{i}(W_{i},\,p^{*}_{\pi})-y(p^{*}_{\pi})\right)\right)S_{i}\right]=0,$
$\mathbb{E}\left[\left\lVert\left(Y_{i}(W_{i},\,p^{*}_{\pi}+h_{n}S_{i})-y(p^{*}_{\pi}+h_{n}S_{i})-\left(Y_{i}(W_{i},\,p^{*}_{\pi})-y(p^{*}_{\pi})\right)\right)S_{i}\right\rVert_{2}^{2}\right]$
$\quad\leq\mathbb{E}\left[\left\lVert Y_{i}(W_{i},\,p^{*}_{\pi}+h_{n}S_{i})-y(p^{*}_{\pi}+h_{n}S_{i})-\left(Y_{i}(W_{i},\,p^{*}_{\pi})-y(p^{*}_{\pi})\right)\right\rVert_{\infty}^{2}\,\big{|}\,S_{i}\right]=\mathcal{O}\left(h_{n}\right).$
These terms are all independent in $i$ so
$\mathbb{E}\left[\left\lVert\frac{1}{n}\sum_{i=1}^{n}\left(Y_{i}(W_{i},\,p^{*}_{\pi}+h_{n}S_{i})-y(p^{*}_{\pi}+h_{n}S_{i})-\left(Y_{i}(W_{i},\,p^{*}_{\pi})-y(p^{*}_{\pi})\right)\right)S_{i}\right\rVert_{2}^{2}\right]=\mathcal{O}\left(h_{n}/n\right),$
and so, by Chebyshev’s inequality and again because $n^{-1}\mathbf{S}^{\top}\mathbf{S}\rightarrow_{p}I_{J\times J}$, the last term in our previous expression for $E_{Y}$ is $\mathcal{O}_{P}(\sqrt{h_{n}/n})$. We have thus shown that
$E_{Y}=h_{n}\xi^{\top}_{y}+\text{vec}\left(Y_{i}(W_{i},\,p^{*}_{\pi})\right)^{\top}\mathbf{S}\left(\mathbf{S}^{\top}\mathbf{S}\right)^{-1}+o_{p}(1/\sqrt{n}),$
and one can derive a similar expression for $E_{Z}$.
@@
$E_{Y}E_{Z}^{-1}=\left(h_{n}\xi^{\top}_{y}+\mathbf{Y}^{*\top}\mathbf{S}\left(\mathbf{S}^{\top}\mathbf{S}\right)^{-1}\right)\left(h_{n}\xi_{z}+\mathbf{Z}^{*\top}\mathbf{S}\left(\mathbf{S}^{\top}\mathbf{S}\right)^{-1}\right)^{-1}+o_{p}\left(\frac{1}{h_{n}\sqrt{n}}\right).$
Now, notice that $\mathbf{Y}^{*}$ and $\mathbf{Z}^{*}$ are independent of $\mathbf{S}$ so
$\mathbf{Y}^{*\top}\mathbf{S}\left(\mathbf{S}^{\top}\mathbf{S}\right)^{-1},\ \ \mathbf{Y}^{*\top}\mathbf{S}\left(\mathbf{S}^{\top}\mathbf{S}\right)^{-1}=\mathcal{O}\left(1/\sqrt{n}\right),$
whereas $h_{n}$ goes to zero slower than $1/\sqrt{n}$ and $\xi_{z}$ is full rank by Assumption 3. Thus, we can Taylor-expand the matrix inverse to get
$E_{Y}E_{Z}^{-1}=\xi^{\top}_{y}\xi_{z}^{-1}+h_{n}^{-1}\mathbf{Y}^{*\top}\mathbf{S}\left(\mathbf{S}^{\top}\mathbf{S}\right)^{-1}\xi_{z}^{-1}-h_{n}^{-1}\xi^{\top}_{y}\xi_{z}^{-1}\mathbf{Z}^{*\top}\mathbf{S}\left(\mathbf{S}^{\top}\mathbf{S}\right)^{-1}\xi_{z}^{-1}+o_{p}\left(\frac{1}{h_{n}\sqrt{n}}\right),$
where the last step is by Taylor expansion. Then, because $n\left(\mathbf{S}^{\top}\mathbf{S}\right)^{-1}\xi_{z}^{-1}\rightarrow_{p}\xi_{z}^{-1}$ and by Slutsky’s lemma,
$E_{Y}E_{Z}^{-1}=\xi^{\top}_{y}\xi_{z}^{-1}+\frac{1}{nh_{n}}\sum_{i=1}^{n}\left(Y_{i}(W_{i},\,p^{*}_{\pi})-\xi_{y}^{\top}\xi_{z}^{-1}Z_{i}(W_{i},\,p^{*}_{\pi})\right)S_{i}^{\top}\xi_{z}^{-1}+o_{p}\left(1/(h_{n}\sqrt{n})\right).$
Finally, recalling the decomposition (61) for $\hat{\tau}_{\text{AIE}}$, the fact that $\tau^{*}_{\text{AIE}}=-\xi^{\top}_{y}\xi_{z}^{-1}\tau^{*,z}_{\text{ADE}}$ and that estimation errors in $\hat{\tau}_{\text{ADE}}^{z}$ are lower order, we see that
$\hat{\tau}_{\text{AIE}}-\tau^{*}_{\text{AIE}}=-\frac{1}{nh_{n}}\sum_{i=1}^{n}\left(Y_{i}(W_{i},\,p^{*}_{\pi})-\xi_{y}^{\top}\xi_{z}^{-1}Z_{i}(W_{i},\,p^{*}_{\pi})\right)S_{i}^{\top}\xi_{z}^{-1}\tau^{*,z}_{\text{ADE}}+o_{p}\left(1/(h_{n}\sqrt{n})\right).$